OpenStudy (anonymous):

Tommy can paint a fence in 13 hours. If his friend Huck helps, they can paint the fence in 1 hour. How fast could Huck paint the fence by himself?

OpenStudy (anonymous):

13 hours to paint the fence that mean in one hour 1/13 of work will be done

OpenStudy (anonymous):

(1/13 + 1/h) t = 1

OpenStudy (anonymous):

if huck paints the fence by himself it would have took him 6 and a half hours

OpenStudy (amistre64):

1/13 + 12/13 = 1 hours work by both. 13/12 hours = huck alone